# LibMoE: A Library for Comprehensive Research Mixture of Experts in Large Language Models

### v2

@@ Abstract
Mixture of experts (MoE) architectures have become a cornerstone for scaling up and are a key component in most large language models such as GPT-OSS, DeepSeek-V3, Llama-4, and Gemini-2.5. However, systematic research on MoE remains severely constrained by the prohibitive computational costs of training and evaluation, restricting large-scale studies accessible to most researchers. We introduce LibMoE, a unified framework for reproducible, efficient, and extensible MoE research that supports both pretraining and sparse-upcycling regimes. Beyond unified implementations, the framework provides transparent analytical tools for probing routing and expert dynamics. Leveraging this foundation, we conduct a comprehensive analysis along three dimensions: (i) routing dynamics, covering expert selection patterns, routing stability and optimality, and how routing entropy reveals task specialization and expert diversity; (ii) the effect of lightweight initialization on load balancing, demonstrating how subtle changes in router initialization shape early expert utilization; and (iii) training regime differences, revealing how sparse upcycling and full pretraining exhibit distinct routing patterns and stability profiles. By lowering the barrier to entry and standardizing evaluation, along with our comprehensive analysis, LibMoE broadens access to MoE research and establishes a reliable benchmark to guide future innovations. Project page: https://fsoft-aic.github.io/fsoft-LibMoE.github.io.
@@ 1 Introduction
However, despite rapid algorithmic advances, progress remains hampered by the lack of standardized frameworks: large-scale MoE research is often accessible only to groups with massive compute resources, while small-scale studies yield fragmented and incomparable results. For instance, training OLMoE-1B-7B required 256 H100 GPUs (Muennighoff et al., 2024), while MOMA experiments utilized 256 A100 GPUs (Lin et al., 2024b). In contrast, most researchers are limited to exploring their ideas in small-scale settings (Nielsen et al., 2025; Han et al., 2024a; Teo & Nguyen, 2024; Csordás et al., 2023; 2024; Nguyen et al., 2025a; Teo & Nguyen, 2025) or even resort to synthetic datasets (Nguyen et al., 2024c; b; Yan et al., 2024). This underscores the pressing need for a transparent and reproducible framework that enables systematic analysis of SMoE behavior with limited constraints, particularly in addressing fundamental design questions such as the trade-off between total and active parameters, the choice between many small versus few large experts, the role of shared experts, scaling dynamics, and the selection of routing algorithms.
@@
To address these issues, we introduce LibMoE, a unified framework that integrates state-of-the-art SMoE algorithms, standardized training pipelines, and transparent analytical tools for studying routing dynamics. In particular, we set two primary goals: (1) To offer a streamlined toolkit to facilitate SMoE research in LLMs within resource-constrained environments; and (2) To perform an in-depth analysis of SMoE routing mechanisms, thereby providing deeper insights into the behaviors. Through this study, we seek to foster and expedite the development of the open-source SMoE community.
Designing LibMoE Framework. To bridge the gap in accessible and reproducible research on MoE, we introduce LibMoE, a unified framework for training and benchmarking SMoE-based models under constrained computational budgets. The framework integrates seven recent state-of-the-art MoE algorithms and supports evaluation across multiple scales and domains, including language modeling with 158M and 679M parameters, as well as vision–language modeling with 5.67B parameters. By providing standardized implementations and fair evaluation protocols, LibMoE lowers the barrier to entry for MoE research while enabling systematic comparisons and scalable experimentation.
@@
Comprehensive Analysis of Mixture-of-Experts Dynamics.Beyond implementation, LibMoE serves as a transparent and extensible platform for probing the internal behaviors of MoE models. It provides modular hooks for tracking routing decisions, expert utilization, load balancing, and inter-expert dynamics, enabling fine-grained inspection of how routing strategies evolve throughout training and inference. Building on this infrastructure, we conduct the first large-scale, cross-domain empirical study of seven recent SMoE algorithms across language modeling (158M, 679M) and vision–language modeling (5.67B). Our analysis uncovers previously underexplored behaviors along three key axes: (i) routing dynamics, revealing differences in expert selection stability, entropy, and specialization; (ii) the overlooked role of lightweight router initialization in shaping early load balancing; and (iii) distinct routing patterns between full pretraining and sparse upcycling regimes. These findings provide a unified and reproducible view into the dynamic nature of SMoE models, offering actionable insights for designing more stable and efficient routing strategies. We believe this study contributes a valuable diagnostic lens to the community, helping both practitioners and researchers better understand and improve the next generation of MoE systems.
@@ 2.1 Mixture of Experts
Originally introduced by Jacobs et al. (Jacobs et al., 1991), the MoE framework was first formulated as an ensemble method that combines multiple specialized models through an adaptive gating network. Early variants, including Jacobs’ gating network and the Hierarchical MoE (HME) (Jordan & Jacobs, 1994), improved learning algorithms for multiclass classification (Chen et al., 1999), and achieved success in domains such as speech recognition (Gales & Airey, 2006). Eigen et al. (2013) extends the MoE to a layer in neural network, which consists of a set of experts (neural networks) and a trainable gate. These early efforts collectively shaped the theoretical and empirical foundation of MoE, setting the stage for the large-scale sparse MoE architectures that gained prominence after 2017.
In 2017, Shazeer et al. (2017b) introduced the SMoE model, which incorporated MoE layers into long short-term memory (LSTM) networks to dramatically scale model capacity without increasing computational cost. Building on this idea, GShard (Lepikhin et al., 2020) extended MoE to Transformers by replacing the feed-forward network (FFN) layers in the T5 (Raffel et al., 2020) architecture with MoE layers. Switch Transformers (Fedus et al., 2022) further simplified the routing mechanism and scaled training to 1.6 trillion parameters, setting a milestone for large language models (LLMs).
Since then, substantial progress has been made in advancing sparsely-gated MoE architectures for LLM development, including novel routing techniques (Chi et al., 2022; Roller et al., 2021; Zuo et al., 2021; Zhong et al., 2024; Muqeeth et al., 2023; Wu et al., 2024b; Wang et al., 2024d; Nielsen et al., 2025; Nguyen et al., 2025c; Csordás et al., 2023), stability-oriented strategies (Zoph et al., 2022; Dai et al., 2022b), fine-grained expert segmentation (Park et al., 2024; Dai et al., 2024; He, 2024), shared experts (Rajbhandari et al., 2022; Liu et al., 2024a; c; Dai et al., 2024; Team, 2024), special-purpose experts (Jin et al., 2024; Yan et al., 2025), gradient update modification (Panda et al., 2025; Yang et al., 2024b), and sparse upcycling (Komatsuzaki et al., 2022; Nakamura et al., 2025; Chen et al., 2025; Hui et al., 2024); with extensions also demonstrated in multimodal learning (Shen et al., 2023; Han et al., 2024a; Li et al., 2024b; Lin et al., 2024a; Wu et al., 2024c; Yun et al., 2024).
@@
Despite the encouraging progress, a clear discrepancy remains in current MoE research. In particular, despite a proliferation of advanced algorithms and theoretical analyses, the most capable and widely deployed LLMs (Agarwal et al., 2025; Comanici et al., 2025; Liu et al., 2024a; c; Dai et al., 2024; xAI, 2024; Yang et al., 2024a; Jiang et al., 2024; Meta AI, 2025; He et al., 2024; Abdin et al., 2024; Databricks, 2024; Wu et al., 2024a) still rely on SMoE variants closely aligned with the original formulation (Shazeer et al., 2017b; Fedus et al., 2022). This indicates that our current theoretical understanding remains largely confined to academic settings and has yet to meaningfully influence real-world deployments precisely the domain where SMoE could be most impactful. We attribute this gap primarily to the high entry barrier posed by massive datasets and computational requirements, which place large-scale SMoE experimentation out of reach for most research groups. Therefore, we develop LibMoE to enable training and evaluation settings that closely reflect real-world practice, including both early-stage pretraining and late-stage sparse upcycling under limited resources. This design allows for fair, comprehensive, and large-scale benchmarking of SMoE algorithms, while making state-of-the-art methods more accessible to the broader research community.
@@ 2.2 Analyzing and Understanding Mixture-of-Experts Models
Since the sparsely-gated MoE layer of Shazeer et al. (2017a), research has explored both its promise of conditional capacity and the challenge of expert imbalance. This motivated stability-focused methods such as StableMoE (Dai et al., 2022a), routing redesigns like Expert-Choice (Zhou et al., 2022), and system optimizations such as MegaBlocks (Gale et al., 2022). Differentiable routing mechanisms, including DSelect-k (Hazimeh et al., 2021), BASE layers (Lewis et al., 2021), and ReMoE (Wang et al., 2024c), further improved flexibility, while AdaMoE (Zeng et al., 2024) and loss-free balancing (Wang et al., 2024a) enhanced efficiency and adaptability.
Beyond algorithmic advances, theoretical and empirical studies examined MoE dynamics. Zhao et al. (2024) derived sparsity-aware generalization bounds, while Fan et al. (2024) ablated routing granularity and expert counts, revealing token-level routers capture syntactic rather than semantic patterns. Nguyen et al. (2024a) characterized cosine routers, and Antoine et al. (2024) showed routers correlate with part-of-speech, highlighting linguistic specialization. Together, these works connect routing design to generalization, specialization, and stability.
Open-source efforts such as OpenMoE (Xue et al., 2024), OLMoE (Muennighoff et al., 2024), and FLAME-MoE (Kang et al., 2025) improved accessibility by releasing models, training frameworks, and diagnostic tools. Despite these contributions, their adoption is often constrained by the need for extensive computational resources, which limits reproducibility and practical experimentation for most researchers. Moreover, these studies are typically tailored to specific design choices, reducing their flexibility to accommodate and compare diverse MoE methods. These efforts about analyzing and understanding mixture of experts models shows the interest as well as the need in analyzing the MoE dynamics.
From a systems perspective, several popular toolkits including Megatron-LM (Shoeybi et al., 2019), FastMoE (He et al., 2021), and Tutel (Hwang et al., 2023) provide implementations of mixture of experts. However, these frameworks present notable limitations for researchers aiming to explore state-of-the-art SMoE variants. For example, Tutel and OpenMoE+t5x primarily target large-scale pre-training across many GPUs, limiting their practicality in resource-constrained environments. Meanwhile, FastMoE offers an outdated implementation that lacks support for recent LLM architectures and advanced distributed training libraries such as DeepSpeed (Lian et al., 2024).
@@ 3.1 Preliminary: Mixture of Experts
The standard SMoE layer (Shazeer et al., 2017b) consists of a router ${\mathcal{R}}(\cdot,W_{r})$, parameterized by $W_{r}$, and a set of $N$ experts $\{g(\cdot,W_{e_{i}})\}_{i=1}^{N}$, each with parameters $W_{e_{i}}$ for $i\in[N]$. Given an input token $\bm{x}$}, the router computes an affinity score vector over all experts as: $vs_{{\mathcal{R}}}=\sigma\left(\mathrm{TopK}_{-\infty}({\bm{x}}^{\top}W_{r})\right)$, where $\sigma$ is a scoring function, typically implemented as a softmax or sigmoid. The operator $\mathrm{TopK}_{-\infty}$ retains the top-$K$ values and sets the remaining entries to negative infinity (-$\infty$), enforcing sparsity. The SMoE output is then calculated as a weighted sum of expert outputs, modulated by the affinity scores:
@@ 3.2 Vision-Language Model: LLaVA Architecture
A key challenge in training SMoE models is obtaining a massive dataset and a large amount of compute. In this work, we propose to incorporate SMoE training into any existing dense LLM checkpoints via the Sparse Upcycling technique (Komatsuzaki et al., 2022), which duplicates the original model to create experts and continue training them on a downstream dataset as a normal SMoE. Consequently, we can bypass the expensive pre-training step and evaluate SMoE algorithms with the most advanced public LLMs.
@@
For training, we adopt the vision-language pre-training task, which is one of the more challenging problem and only requires a rather small amount of data to start (around 1e9 tokens). To this end, we follow the CUMO framework (Li et al., 2024b) to upcycle the LLaVA model (Liu et al., 2023a), which consists of three modules: a pre-trained visual encoder, a pre-trained LLM, and a randomly initialized MLP connector. Training follows a two-stage process. First, the dense training stage initializes the MLP connector and train it to connect the pre-trained visual encoder to the pre-trained LLM. Second, the SMoE training stage upcyles the model to become SMoE and also trains all components to obtain the visual instruction following capabilities, Figure 1 shows the model training process in detail. Importantly, we follow Li et al. (2024b) to only upcyle the MLP connector and the visual encoder since upcycling a dense LLM is found to be worse than just using an SMoE LLM. Moreover, we highlight that the dense training stage is unrelated to SMoE, and thus the checkpoints can be re-used to train different SMoE algorithms. Overall, in our small scale setting with 4xH100 GPUs, the whole training process takes a bit over 55 hours while the third training stage takes 24 hours, which is an affordable cost.

### v3

@@ Abstract
Mixture of experts (MoE) architectures have become a cornerstone for scaling up and are a key component in most large language models such as GPT-OSS, DeepSeek-V3, Llama-4, and Gemini-2.5. However, systematic research on MoE remains severely constrained by the prohibitive computational costs of training and evaluation, restricting large-scale studies accessible to most researchers. We introduce LibMoE, a unified framework for reproducible, efficient, and extensible MoE research that supports both pretraining and sparse-upcycling regimes. Beyond unified implementations, the framework provides transparent analytical tools for probing routing and expert dynamics. Leveraging this foundation, we conduct a comprehensive analysis along three dimensions: (i) routing dynamics, covering expert selection patterns, routing stability and optimality, and how routing entropy reveals task specialization and expert diversity; (ii) the effect of lightweight initialization on load balancing, demonstrating how subtle changes in router initialization shape early expert utilization; and (iii) training regime differences, revealing how sparse upcycling and full pretraining exhibit distinct routing patterns and stability profiles. By lowering the barrier to entry and standardizing evaluation, along with our comprehensive analysis, LibMoE broadens access to MoE research and establishes a reliable benchmark to guide future innovations. GitHub: https://github.com/Fsoft-AIC/LibMoE.
@@ 1 Introduction
However, despite rapid algorithmic advances, progress in MoE research remains fundamentally hindered by the lack of standardized frameworks. Large-scale MoE experiments are typically feasible only for groups with access to massive computational resources training OLMoE-1B-7B, for example, required 256 H100 GPUs (Muennighoff et al., 2024), and MOMA relied on 256 A100 GPUs (Lin et al., 2024b). In contrast, the majority of researchers are confined to small-scale studies (Nielsen et al., 2025; Han et al., 2024a; Teo and Nguyen, 2024; Csordás et al., 2023; 2024; Nguyen et al., 2025a; Teo and Nguyen, 2025) or even synthetic benchmarks (Nguyen et al., 2024c; b; Yan et al., 2024), resulting in fragmented and often incomparable results. This persistent discrepancy severely undermines the potential of MoE architectures, whose true strengths emerge only in large-scale training regimes. As a result, even the most prominent deployed models such as Phi3 (Abdin and others, 2024), Skywork-MoE (Wei et al., 2024), and OSS-GPT (OpenAI, 2025) tend to rely on conventional SMoE designs (Fedus et al., 2022), with recent advances in routing and expert selection algorithms seeing little adoption in practice (Do et al., 2023; Zhou et al., 2022; Dai et al., 2022b; Wang et al., 2024d). These challenges underscore a critical need for a unified, reproducible framework that lowers the barrier to meaningful SMoE research under realistic resource constraints and enables rigorous, systematic evaluation of both established and emerging methods.
@@
To address these challenges, we present LibMoE a unified framework that brings together state-of-the-art SMoE algorithms, standardized training pipelines, and transparent analytical tools for rigorous investigation of routing dynamics. Our primary objectives are twofold: (1) to provide an accessible and streamlined toolkit that enables meaningful SMoE research on large language models (LLMs) under realistic resource constraints; and (2) to enable comprehensive, reproducible analysis of SMoE routing mechanisms, yielding deeper insights into expert behaviors. With LibMoE, we aim to accelerate progress and foster a more collaborative, open-source SMoE research community.
Designing the LibMoE Framework. To address persistent challenges in accessible and reproducible MoE research, we introduce LibMoE a unified framework for systematic evaluation of SMoE methods across both full pretraining and sparse-upcycling regimes (Komatsuzaki et al., 2022). LibMoE accommodates a broad spectrum of model sizes and algorithmic variants, supporting rigorous comparisons from small-scale language models (0.15B and 0.68B parameters) to large-scale vision–language models (5.67B parameters). The framework integrates seven recent state-of-the-art SMoE algorithms and provides standardized pipelines for both pretraining and evaluation, enabling consistent benchmarking and cross-domain analysis. All MoE algorithm experiments are conducted under realistic resource constraints, with training times ranging from 6 to 44 hours on $4\times$ H100 GPUs (see Table 11 for details). While our results reveal that current SMoE algorithms yield only marginal performance improvements under these settings, LibMoE substantially lowers the barrier to entry, making advanced and reproducible experimentation feasible even for research groups with limited computational resources.
@@
Comprehensive Analysis of Mixture-of-Experts Dynamics. Beyond providing unified implementations, LibMoE is designed as a transparent and extensible infrastructure for systematically analyzing the behavior of SMoE models. The framework exposes modular instrumentation for monitoring routing decisions, expert utilization, load balancing, and inter-expert interactions, enabling fine-grained inspection of routing behavior throughout both training and inference. Using this analytical capability, we conduct a large-scale, cross-domain empirical study of seven recent SMoE algorithms across multiple training regimes, including small-scale pretraining (0.15B and 679M parameters) and sparse upcycling at the 5.67B scale (Komatsuzaki et al., 2022). Our analysis uncovers consistent and previously under-characterized differences along three dimensions: (i) routing dynamics, reflected in expert selection stability, entropy evolution, and specialization patterns; (ii) the impact of lightweight router initialization on early-stage load balancing; and (iii) systematic divergences in routing behavior between full pretraining and sparse upcycling regimes. Together, these results demonstrate how design choices in routing and initialization influence training stability and expert utilization, and establish a reproducible analytical basis for principled comparison of SMoE algorithms.
@@ 2.1 Mixture of Experts
Originally introduced by Jacobs et al. (1991), the MoE framework was first formulated as an ensemble method that combines multiple specialized models through an adaptive gating network. Early variants, including Jacobs’ gating network and the Hierarchical MoE (HME) (Jordan and Jacobs, 1994), improved learning algorithms for multiclass classification (Chen et al., 1999), and achieved success in domains such as speech recognition (Gales and Airey, 2006). Eigen et al. (2013) extends the MoE to a layer in neural network, which consists of a set of experts (neural networks) and a trainable gate. These early efforts collectively shaped the theoretical and empirical foundation of MoE, setting the stage for the large-scale sparse MoE architectures that gained prominence after 2017. In 2017, Shazeer et al. (2017b) introduced the SMoE model, which incorporated MoE layers into long short-term memory (LSTM) networks to dramatically scale model capacity without increasing computational cost. Building on this idea, GShard (Lepikhin et al., 2020) extended MoE to Transformers by replacing the feed-forward network (FFN) layers in the T5 (Raffel et al., 2020) architecture with MoE layers. Switch Transformers (Fedus et al., 2022) further simplified the routing mechanism and scaled training to 1.6 trillion parameters, setting a milestone for large language models (LLMs). Since then, substantial progress has been made in advancing sparsely-gated MoE architectures for LLM development, including novel routing techniques (Chi et al., 2022; Roller et al., 2021; Zuo et al., 2021; Zhong et al., 2024; Muqeeth et al., 2023; Wu et al., 2024b; Wang et al., 2024d; Nielsen et al., 2025; Nguyen et al., 2025c; Csordás et al., 2023), stability-oriented strategies (Zoph et al., 2022; Dai et al., 2022b), fine-grained expert segmentation (Park et al., 2024; Dai et al., 2024; He, 2024), shared experts (Rajbhandari et al., 2022; Liu et al., 2024a; c; Dai et al., 2024; Team, 2024), special-purpose experts (Jin et al., 2024; Yan et al., 2025), gradient update modification (Panda et al., 2025; Yang et al., 2024b), and sparse upcycling (Komatsuzaki et al., 2022; Nakamura et al., 2025; Chen et al., 2025; Hui et al., 2024); with extensions also demonstrated in multimodal learning (Shen et al., 2023; Han et al., 2024a; Li et al., 2024b; Lin et al., 2024a; Wu et al., 2024c; Yun et al., 2024). Despite the encouraging progress, a clear discrepancy remains in current MoE research. In particular, despite a proliferation of advanced algorithms and theoretical analyses, the most capable and widely deployed LLMs (Agarwal et al., 2025; Comanici et al., 2025; Liu et al., 2024a; c; Dai et al., 2024; xAI, 2024; Yang et al., 2024a; Jiang et al., 2024; Meta AI, 2025; He et al., 2024; Abdin and others, 2024; Databricks, 2024; Wu et al., 2024a) still rely on SMoE variants closely aligned with the original formulation (Shazeer et al., 2017b; Fedus et al., 2022). This indicates that our current theoretical understanding remains largely confined to academic settings and has yet to meaningfully influence real-world deployments precisely the domain where SMoE could be most impactful. We attribute this gap primarily to the high entry barrier posed by massive datasets and computational requirements, which place large-scale SMoE experimentation out of reach for most research groups. Therefore, we develop LibMoE to enable training and evaluation settings that closely reflect real-world practice, including both early-stage pretraining and late-stage sparse upcycling under limited resources. This design allows for fair, comprehensive, and large-scale benchmarking of SMoE algorithms, while making state-of-the-art methods more accessible to the broader research community.
@@ 2.2 Analyzing and Understanding Mixture-of-Experts Models
Since the introduction of the sparsely-gated MoE layer (Shazeer et al., 2017a), research has explored both the promise of conditional capacity and the challenge of expert imbalance. This has driven the development of stability-oriented methods (e.g., StableMoE (Dai et al., 2022a)), novel routing designs (e.g., Expert-Choice (Zhou et al., 2022)), and system-level optimizations (e.g., MegaBlocks (Gale et al., 2022)). Differentiable routing mechanisms, including DSelect-k (Hazimeh et al., 2021), BASE layers (Lewis et al., 2021), and ReMoE (Wang et al., 2024c), have further increased flexibility, while AdaMoE (Zeng et al., 2024) and loss-free balancing (Wang et al., 2024a) have enhanced efficiency and adaptability. Beyond algorithmic innovation, a growing body of theoretical and empirical work has examined MoE dynamics. For example, Zhao et al. (2024) derived sparsity-aware generalization bounds, while Fan et al. (2024) systematically ablated routing granularity and expert count, revealing that token-level routers tend to capture syntactic rather than semantic patterns. Other studies have characterized router behavior (e.g., Nguyen et al. (2024a)) and highlighted linguistic specialization (e.g., Antoine et al. (2024)), collectively connecting routing design to generalization, specialization, and stability. Efforts to democratize MoE research have yielded open-source frameworks such as OpenMoE (Xue et al., 2024), OLMoE (Muennighoff et al., 2024), and FLAME-MoE (Kang et al., 2025), which provide models, training frameworks, and diagnostic tools. However, the adoption of these resources is often limited by substantial computational requirements, restricting reproducibility and practical experimentation. Moreover, most prior studies focus on behaviors specific to their own MoE variants, rather than providing systematic cross-method comparisons. In contrast, our work aims to provide a unified and comprehensive empirical analysis of leading MoE algorithms under a standardized evaluation protocol. By systematically comparing a diverse set of methods and probing the effects of key design parameters, we offer deeper insights into the factors driving performance and specialization in modern MoE architectures.
2.3 Mixture of Experts Toolkits
Several open-source toolkits, including FastMoE (He et al., 2021), OpenMoE+t5x (Xue et al., 2024), and Tutel (Hwang et al., 2023), support the implementation of SMoE algorithms. However, these frameworks present notable limitations for contemporary research. Tutel and OpenMoE+t5x are primarily designed for large-scale pretraining on hundreds of GPUs, restricting accessibility for groups with limited resources. FastMoE, while more lightweight, lacks support for recent LLM architectures and advanced distributed training libraries such as DeepSpeed (Lian et al., 2024). In contrast, LibMoE is specifically designed to lower these barriers: researchers can train models with as few as 11B tokens for vision–language tasks and 66B tokens for language modeling representing a $1{,}000\times$ reduction in data requirements compared to OpenMoE while benefiting from support for modern LLMs and modular distributed training.
@@ 3.1 Preliminary: Mixture of Experts
The standard SMoE layer (Shazeer et al., 2017b) consists of a router ${\mathcal{R}}(\cdot,W_{r})$, parameterized by $W_{r}$, and a set of $N$ experts $\{g(\cdot,W_{e_{i}})\}_{i=1}^{N}$, each with parameters $W_{e_{i}}$ for $i\in[N]$. Given an input token $\bm{x}$}, the router computes an affinity score vector over all experts as: ${\bm{s}}_{{\mathcal{R}}}=\sigma\left(\mathrm{TopK}_{-\infty}({\bm{x}}^{\top}W_{r})\right)$, where $\sigma$ is a scoring function, typically implemented as a softmax or sigmoid. The operator $\mathrm{TopK}_{-\infty}$ retains the top-$K$ values and sets the remaining entries to negative infinity (-$\infty$), enforcing sparsity. The SMoE output is then calculated as a weighted sum of expert outputs, modulated by the affinity scores:
@@ 3.2 Vision-Language Model: LLaVA Architecture
A key challenge in training SMoE models is obtaining a massive dataset and a large amount of compute. Thus, beyond the traditional pre-training setting, we propose to incorporate SMoE training into any existing dense LLM checkpoints via the Sparse Upcycling technique (Komatsuzaki et al., 2022), which duplicates the original model to create experts and continue training them on a downstream dataset as a normal SMoE. Consequently, we can bypass the expensive pre-training step and evaluate SMoE algorithms with the most advanced public LLMs.
@@
[…]